OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 7 and a9 = 55 ?

OpenStudy (amistre64):

an arith seq is defined as: an = a1 + d(n-1) to determine "d" an - a1 ------ = d n-1

OpenStudy (anonymous):

A) 151 B) 176 C) 199 D) 224

OpenStudy (anonymous):

d = 6

hartnn (hartnn):

yes,very good,d indeed equals 6 now to find a25,just put a1=7,d=6,n=25 in the an equation given in 1st comment....

OpenStudy (anonymous):

an = a1 + d(n-1) .... so 25 = 7 + 6(24) ?

OpenStudy (amistre64):

perfect

OpenStudy (amistre64):

well, a{25} is the name of the 25th term

hartnn (hartnn):

an = 7 + 6(24) = 25th term

OpenStudy (amistre64):

the value of a{25} = 7 + 6(24)

OpenStudy (anonymous):

151 ?

hartnn (hartnn):

yes,u are correct :) 7+144=151

OpenStudy (anonymous):

great thanks :)